The head of a tropical fish is 1/3 as long as its midsection. Its tail is as long as its head and midsection combined. The total length of the fish is 48cm. How long is the tail of the fish?
Answer by jermz69(5)

The head of a tropical fish is 1/3 as long as its midsection. Its tail is as long as its head and midsection combined. The total length of the fish is 48cm. How long is the tail of the fish?

Let x: midsection
1/3x: head
x + 1/3x: tail

Equation:
x + 1/3x + x + 1/3x = 48
8/3x = 48
3(8/3x) = 3(48)
8x = 144
x = 24

tail = x + 1/3x
24 + 1/3(24)
24 + 8
32
tail = 32

I'm jermz69. Hope I helped!